LASER SPECTROSCOPY OF BaOH AND BaOD: ANOMALOUS SPIN-ORBIT COUPLING IN THE Ã^{2}Π STATE

\underline{J. D. TANDY}, J.-G. WANG, and P. F. BERNATH, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK.

The Ã^{2}Π − X̃^{2}Σ^{+} transition of BaOD has been rotationally analyzed using high-resolution laser excitation spectroscopy. BaOD molecules were synthesized in a Broida-type oven and detected using a single mode Ti:Sapphire laser. Measured rotational lines have been assigned, and rotational and fine structure parameters determined through a combined leastsquares fit with the millimeter-wave pure rotational data of the X̃^{2}Σ^{+} state. A significantly different spin-orbit coupling constant from the corresponding value for BaOH was observed and attributed to global and local perturbations arising from vibrationally excited bands of the Ã^{′2}Δ state. Λ-doubling constants for the Ã^{2}Π state also showed poor agreement with the predictions of the pure precession model. To further understand the nature of the interactions between the B̃^{2}Σ^{+}, Ã^{2}Π and Ã^{′2}Δ states of BaOH, a V-type optical-optical double resonance spectroscopy experiment was prepared to locate the lower-lying excited state levels. Preliminary results have yielded a band in close proximity on the lower wavenumber side of the Ã^{2}Π_{1/2} spin component of BaOH with a relatively large p value. Further results of this ongoing experiment will be presented.